# Evolutionary algorithms meet self-supervised learning: a comprehensive survey

### v1

@@ IV-A3 Pretext task – Learning
For example, Cheng et al. [69] evolve hyperparameters for sparse AEs. Their approach does not enforce the same sparsity penalty coefficient for all neurons. Instead, each neuron can have its own coefficient. The authors propose an EA that evolves both the set of individual penalty coefficients and the learning rate . The evolutionary process is guided by two objectives derived from the loss function: one minimises the reconstruction loss, and the other minimises the sparsity penalty term. The method is evaluated on multiple network architectures using an image classification downstream task. Additionally, an ablation study is conducted to assess the impact of the number of labelled samples used in downstream training.
@@
In eVAE, a coefficient is introduced to control the relative importance of the KL divergence term, and this value is evolved. The evolutionary process operates by first training a variational AE for epochs. Then, a population of candidate values is generated. These are evaluated based on the direction of the stochastic gradient and are varied using variational crossover and mutation. The best value is selected and returned to eVAE, initiating another round of training.
@@
EvolQ [73] is a post-training quantisation algorithm that compresses the weights of a vision transformer, guided by a contrastive SSL loss. The first step of EvolQ involves learning the scale factors applied to weight vectors and activations. The evolutionary process is applied separately to each transformer block. Within each block, the EA generates a population of perturbation vectors that modify the learned scales. These vectors are evolved to find the quantised configuration that minimises the contrastive loss.
@@
IV-A4 Pretext task – Multiple components
@@
Charte et al. [81, 82] explore the use of several EAs to evolve AEs. Their approach uses a fixed-length genotype, where genes encode topological features such as the type of AE, the number of layers, and the size and activation function of each layer, assuming a symmetric topology. The final gene specifies the loss function used during training. The fitness function combines training loss, the number of layers, and the dimensionality of the learned representations. The authors evaluate this setup using GA, ES, and Differential Evolution (DE) algorithms.
@@
Silhan et al.[85] propose a layer-wise evolutionary strategy for AEs, where a new layer is added to both the encoder and decoder at each step to maintain symmetry. The EA evolves two topological parameters – dropout ratio and activation function – and two learning parameters – learning rate and momentum. At each generation, these hyperparameters are mutated, and weights are inherited by offspring, following a Lamarckian evolution strategy. Each individual is evaluated using the Local Continuity Meta-Criterion (LCMC)[86], a metric that measures the quality of dimensionality reduction.
@@ IV-A5 Downstream task
Zhou et al. [90] fine-tune a self-supervised pretrained model for a text matching task using supervised contrastive learning. The core idea is to ensure that pairs of matching texts are represented similarly in the latent space, while non-matching pairs are pushed apart. Based on the misclassified examples from the trained model, an EA algorithm is applied to evolve alternative options that match the original text. These new pairs are added to the original set, and another round of contrastive training is performed.
@@
Instead of encoding parameterised topology aspects in the evolutionary process, EAs can be used directly to solve the downstream task. For example, Rodrigues et al. [97] apply evolved individuals on top of a set of representations trained in a self-supervised manner. They use GP to evolve models that solve downstream tasks in the tabular data domain. The authors note that while SSL works seamlessly for structured domains such as NLP and computer vision, it is less effective for tabular data due to its heterogeneity and lack of global structure. Indeed, they show that incorporating SSL representations actually hinders the performance of downstream models.
@@
IV-B SSL for EC
As previously discussed in Section III, SSL models can be integrated into EC by targeting any of the components of an evolutionary algorithm to enhance aspects of the evolution process. This area of research is rapidly expanding, with several works focusing on three key components of an evolutionary algorithm: evolutionary representation, variation operators and fitness evaluation.
IV-B1 Evolutionary Representation
@@
One advantage of this approach is related to the dimensionality of the problem. Latent spaces can be viewed as compressed versions of the input space. By performing evolution in the latent space, the dimensionality of the problem is reduced, making it easier to navigate the search space. For example, Tian et al. [47] use a denoising AE integrated into an EA designed for large-scale multiobjective optimisation problems. Before each generation, an AE is trained using non-dominated individuals, which are corrupted by randomly setting some chromosome values to zero. The AE encodes solutions into the latent space, where genetic operators are applied. The solutions are then decoded back into the original search space for evaluation.
Cui et al. [99] frame an EA as a cooperative co-evolution problem with two populations to balance exploration and exploitation. One population performs evolution in the original search space, representing exploitation, while the other undergoes dimensionality reduction through an AE and applies genetic operators in the reduced search space. The populations are split at the start of each generation, with the individuals with the worst fitness being assigned to the population fed to the AE. This population is classified as performing exploration. In the initial generations, evolution does not rely on the AE since it is gathering individuals for training.
@@ IV-B3 Fitness
V Challenges and future research directions
E-SSL addresses challenges in representation learning and algorithmic optimisation. The intersection of EC and SSL can be beneficial to automate the design of neural networks when the labelled data is scarce [88], or when there is a need to design networks that can be transferred to multiple tasks [64]. Additionally, E-SSL can be used to improve different aspects of EA algorithms without depending on labels. Nevertheless, five main challenges were identified in this field, aiming to provide guidance on future work to be done within E-SSL.
V-A Pretext task design automation
The most noticeable trend in E-SSL is the preference for auto-encoding models, along with a lack of exploration into multiple branch networks. Although AE models can be competitive, what these competitive proposals share is the incorporation of a masking objective instead of the original reconstruction-based loss [31, 120]. This underscores the importance of researching the pretext task that produces the best possible representation. The works surveyed in this context approach pretext task learning as a tool to maintain model performance with more limited resources, optimise coefficients of a loss function associated with a single pretext task, or define a combinatorial optimisation problem by framing pretext task learning as a linear combination of multiple off-the-shelf pretext tasks. However, one possible avenue for future research could be exploring the evolution of new pretext tasks that could facilitate representation learning.
The design of pretext tasks, however, depends on the architecture chosen for the first stage. More specifically, SSL architectures rely on an encoder component that is reused for the downstream task. Other components serve auxiliary roles only during the pretext stage, and their existence depends on the pretext task. In its entirety, the actual model architecture (encoder) is only a small part of a larger structure, which we will refer to as the meta-architecture from hereon. We propose that E-SSL could play a pivotal role in designing meta-architectures for pretext task solutions by providing a unified view. Siamese networks, multiple-branch networks, and AEs can be seen as meta-architectures derived from a common definition but instantiated with different parameter sets. Therefore, it is possible to define a search space of meta-architectures that shapes how an EA explores candidate solutions. The design of pretext tasks under the meta-architecture space is naturally more intricate to define and search, but it has teh potential of producing better representations.
V-B Experiment design under label scarcity scenarios
@@
V-C Exploration of fitness metrics that balance the trade-off between time consumption and fidelity
@@
V-D Fitness evaluation speed-up
@@
V-E Exploration of other meta-architectures
As it was mentioned in earlier challenges, AEs are overall the most commonly adopted meta-architecture. When SSL is applied for EC, this trend is more clear. One exception to this rule is observed in the case when SSL is applied to fitness evaluation aspects, as we did not find instances of AEs being used to train surrogate models for evolutionary NAS fitness assignments. This is in contrast to the operators component, where no variation operators using meta-architectures beyond AEs were observed, nor did we find any E-SSL works designing crossover operators. This highlights potential gaps in the current research that could be explored in future work.
@@ V-E Exploration of other meta-architectures
VI Conclusion
In this survey, we provide a comprehensive review of the intersection between EML and SSL, introducing the emerging field of Evolutionary Self-Supervised learning (E-SSL). We categorise the existing literature into two main groups: one that explores the application of EC to SSL, and another that applies SSL to enhance EC processes. Within these groups, we further refine the classification based on the SSL stage and the specific components of Deep Neural Networks or evolutionary algorithms targeted by SSL.
A clear trend emerges from this review: AE models are dominant across both groups, suggesting their significant role in the development of E-SSL. As this field evolves, there is great potential for E-SSL to uncover novel techniques in representation learning and optimise evolutionary processes. By providing a structured overview of the field, highlighting its challenges, and identifying promising research directions, we aim to catalyse further exploration and development of E-SSL as a dynamic and interdisciplinary area of study.

### v2

@@ 4.1.3 Pretext task – Learning
For example, Cheng et al. [78] evolve hyperparameters for sparse AEs. Their approach does not enforce the same sparsity penalty coefficient for all neurons. Instead, each neuron can have its own coefficient. The authors propose an EC algorithm that evolves both the set of individual penalty coefficients and the learning rate . The evolutionary process is guided by two objectives derived from the loss function: one minimises the reconstruction loss, and the other minimises the sparsity penalty term. The method is evaluated on multiple network architectures using an image classification downstream task. Additionally, an ablation study is conducted to assess the impact of the number of labelled samples used in downstream training.
@@
In eVAE, a coefficient is introduced to control the relative importance of the KL divergence term, and this value is evolved. The evolutionary process operates by first training a variational AE for epochs. Then, a population of candidate values is generated. These are evaluated based on the direction of the stochastic gradient and are changed using ‘variational’ crossover and mutation. The best value is selected and returned to eVAE, initiating another round of training.
@@
EvolQ [74] is a post-training quantisation algorithm that compresses the weights of a vision transformer, guided by a contrastive SSL loss. The first step of EvolQ involves learning the scale factors applied to weight vectors and activations. The evolutionary process is applied separately to each transformer block. Within each block, the evolutionary process generates a population of perturbation vectors that modify the learned scales. These vectors are evolved to find the quantised configuration that minimises the contrastive loss.
@@
4.1.4 Pretext task – Multiple components
@@
Charte et al. [83, 84] explore the use of several EC algorithms to evolve AEs. Their approach uses a fixed-length genotype, where genes encode topological features such as the type of AE, the number of layers, and the size and activation function of each layer, assuming a symmetric topology. The final gene specifies the loss function used during training. The fitness function combines training loss, the number of layers, and the dimensionality of the learned representations. The authors evaluate this setup using GA, ES, and Differential Evolution (DE) algorithms.
@@
Silhan et al. [97] propose a layer-wise evolutionary strategy for AEs, where a new layer is added to both the encoder and decoder at each step to maintain symmetry. The evolutionary process evolves two topological parameters – dropout ratio and activation function – and two learning parameters – learning rate and momentum. At each generation, these hyperparameters are mutated, and weights are inherited by offspring, following a Lamarckian evolution strategy. Each individual is evaluated using the Local Continuity Meta-Criterion (LCMC) [99], a metric that measures the quality of dimensionality reduction.
@@ 4.1.5 Downstream task
Zhou et al. [100] fine-tune a self-supervised pretrained model for a text matching task using supervised contrastive learning. The core idea is to ensure that pairs of matching texts are represented similarly in the latent space, while non-matching pairs are pushed apart. Based on the misclassified examples from the trained model, an EC algorithm is applied to evolve alternative options that match the original text. These new pairs are added to the original set, and another round of contrastive training is performed.
@@
Instead of encoding parameterised topology aspects in the evolutionary process, EC can be used directly to solve the downstream task. For example, Rodrigues et al. [61] apply evolved individuals on top of a set of representations trained in a self-supervised manner. They use GP to evolve models that solve downstream tasks in the tabular data domain. The authors note that while SSL works seamlessly for structured domains such as NLP and computer vision, it is less effective for tabular data due to its heterogeneity and lack of global structure. Indeed, they show that incorporating SSL representations actually hinders the performance of downstream models.
@@
4.2 SSL for EC
As previously discussed in Section 3, SSL models can be integrated into EC by targeting any of the components of an evolutionary algorithm to enhance aspects of the evolution process. This area of research is rapidly expanding, with several works focusing on three key components of an evolutionary algorithm: evolutionary representation, variation operators and fitness evaluation. The studies that fall under this category are described in Table 2. It describes what does a certain work evolve, if it uses an off-the-shelf model in the evolution process or if tries to search new models, the type of goal defined in the pretext task, what part of the model is integrated in the evolutionary process, and what EC component the model targets.
| SSL engagement with EC | |||||
| Work name | What is evolved | Off-the-shelf model | SSL goal type | What is used for evolution | EC component targeted |
| Wei et al. [102] | Neural networks | No | Regression + Contrastive (separately) | Downstream output | Fitness evaluation |
| CLIP MasterPrints [103] | Image representations | Yes | Multimodal agreement | Pretext output | Fitness evaluation |
| EvoProteus [104] | Image generation parameters | Yes | Multimodal agreement | Pretext output | Fitness evaluation |
| Machado et al. [105] | Prompts | Yes | Multimodal agreement | Pretext output | Fitness evaluation |
| BPT-VLM [106] | Prompts | Yes | Multimodal agreement | Pretext output | Fitness evaluation |
| HEP-ENAS [107] | Neural networks | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| Hu et al. [108] | Several optimisation problems | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| DAE-GP [109, 110, 111, 112] | Several optimisation problems | No | Input reconstruction | Pretext output | Operators |
| EAEPSO [113] | Neural networks | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| MOEA/PSL [50] | Several optimisation problems | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| TTLBO [114] | Several optimisation problems | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| ENAO [115] | Neural networks | No | Input reconstruction | Representations | Genotype Phenotype mapping |
| Gong et al. [116] | Neural networks | No | Input reconstruction | Representations | Genotype-Phenotype mapping |
| Shem-Tov et al. [117] | Regression coefficients | No | Masking | Pretext task | Operators |
| DAGA [118] | Several optimisation problems | No | Input reconstruction | Pretext output | Operators |
| DAE-EDA [119] | Several optimisation problems | No | Input reconstruction | Pretext output | Operators |
| Thakkar et al. [120] | Several optimisation problems | No | Input reconstruction | Pretext output | Operators |
| Dang et al. [121] | Neural networks | No | Input reconstruction | Downstream output | Fitness evolution |
4.2.1 Evolutionary Representation
@@
One advantage of this approach is related to the dimensionality of the problem. Latent spaces can be viewed as compressed versions of the input space. By performing evolution in the latent space, the dimensionality of the problem is reduced, making it easier to navigate the search space. For example, Tian et al. [50] use a denoising AE integrated into an evolutionary process designed for large-scale multiobjective optimisation problems. Before each generation, an AE is trained using non-dominated individuals, which are corrupted by randomly setting some chromosome values to zero. The AE encodes solutions into the latent space, where genetic operators are applied. The solutions are then decoded back into the original search space for evaluation.
Cui et al. [114] frame an EC algorithm as a cooperative co-evolution problem with two populations to balance exploration and exploitation. One population performs evolution in the original search space, representing exploitation, while the other undergoes dimensionality reduction through an AE and applies genetic operators in the reduced search space. The populations are split at the start of each generation, with the individuals with the worst fitness being assigned to the population fed to the AE. This population is classified as performing exploration. In the initial generations, evolution does not rely on the AE since it is gathering individuals for training.
@@ 4.2.3 Fitness
5 Challenges and future research directions
E-SSL addresses challenges in representation learning and algorithmic optimisation. The intersection of SSL and EML can be beneficial to automate the design of neural networks when the labelled data is scarce [52], or when there is a need to design networks that can be transferred to multiple tasks [51]. Additionally, E-SSL can be used to improve different aspects of EC algorithms without depending on labels. Nevertheless, five main challenges were identified in this field, aiming to provide guidance on future work to be done within E-SSL.
5.1 Pretext task design automation
The most noticeable trend in E-SSL is the preference for auto-encoding models that rely goals based on input reconstruction, along with a lack of exploration into multiple branch networks. Although AE models can be competitive, what these competitive proposals share is the incorporation of a masking objective instead of the original reconstruction-based loss [34, 123]. This underscores the importance of researching the pretext task that produces the best possible representation. The works surveyed in this context approach pretext task learning as a tool to maintain model performance with more limited resources, optimise coefficients of a loss function associated with a single pretext task, or define a combinatorial optimisation problem by framing pretext task learning as a linear combination of multiple off-the-shelf pretext tasks. However, one possible avenue for future research could be exploring the evolution of new pretext tasks that could facilitate representation learning.
The design of pretext tasks, however, depends on the architecture chosen for the first stage. More specifically, SSL architectures rely on an encoder component that is reused for the downstream task. Other components serve auxiliary roles only during the pretext stage, and their existence depends on the pretext task. In its entirety, the actual model architecture (encoder) is only a small part of a larger structure, which we will refer to as the meta-architecture from hereon. We propose that E-SSL could play a pivotal role in designing meta-architectures for pretext task solutions by providing a unified view. Siamese networks, multiple-branch networks, and AEs can be seen as meta-architectures derived from a common definition but instantiated with different parameter sets. Therefore, it is possible to define a search space of meta-architectures that shapes how EC algorithms explore candidate solutions. The design of pretext tasks under the meta-architecture space is naturally more intricate to define and search, but it has the potential of producing better representations.
5.2 Experiment design under label scarcity scenarios
@@
5.3 Exploration of fitness metrics that balance the trade-off between time consumption and fidelity
@@
5.4 Fitness evaluation speed-up
@@
5.5 Exploration of other meta-architectures
As it was mentioned in earlier challenges, AEs that are trained based on an input reconstruction objective are overall the most commonly adopted meta-architecture. When SSL is applied for EC, this trend is more clear. One exception to this rule is observed in the case when SSL is applied to fitness evaluation aspects, as we did not find instances of AEs being used to train surrogate models for evolutionary NAS fitness assignments. This is in contrast to the operators component, where no variation operators using meta-architectures beyond AEs were observed, nor did we find any E-SSL works designing crossover operators. This highlights potential gaps in the current research that could be explored in future work.
@@ 5.5 Exploration of other meta-architectures
5.6 Summary
The field of E-SSL has more research developed in cases whereby EC is applied for SSL. This is partially fuelled by the amount of studies that apply NAS approaches focused of self-supervised objectives. There is a need to explore alternative objectives and, more importantly, to engage EC directly with the goal in the pretext task via loss function or the meta-architecture designed to train it. This highlights the challenge of using EC to discover alternative forms of representation learning.
There is also a preference for AE models that rely on an input reconstruction objective. From the perspective of applying EC to SSL, this raises the same aforementioned challenge. However, from the point view of applying SSL for EC, it also highlights that there can be alternative components designed with the help of SSL models, which could improve aspects of EC.
Compared to supervised learning, using a model trained with SSL comes with an overhead - it is a 2-step process whose pretext task can be time consuming to train. This means that if one relies on the downstream task to understand the quality of the representations, significant time costs will be incurred on any system that relies on this metric. Therefore, it is important to find reliable and computationally efficient proxy measures that can accurately reflect the quality of learned representations.
When it comes to SSL, the quality of any suggested approach should be measured against different amounts of labelled data, in order to prove its robustness to less labelled data. This consideration should be taken into account not only during the evolutionary process that comes up with a model, but also during any post-evolution methods that are applied to enhance the model obtained by the evolutionary system.
6 Conclusions
This survey provides a comprehensive review of the intersection between EML and SSL, introducing the emerging field of Evolutionary Self-Supervised Learning (E-SSL). We categorise the literature into two groups: (i) studies that apply EC to SSL, and (ii) studies that use SSL to enhance EC. Within each group, we further classify works according to the SSL stage and the specific components of Deep Neural Networks or evolutionary algorithms targeted by SSL.
Our analysis reveals two clear trends: first, most research applies EC to SSL, while the reverse direction remains underexplored; second, AEs trained with input reconstruction objectives dominate current approaches. This indicates both their importance and the opportunity to investigate alternative or evolved SSL objectives.
As this field develops, there is great potential for E-SSL to uncover novel techniques in representation learning and optimise evolutionary processes. By providing a structured overview of the field, highlighting its challenges, and identifying promising research directions, we aim to catalyse further exploration and development of E-SSL as a dynamic and interdisciplinary area of study.
Acknowledgments
This work is financed through national funds by FCT - Fundação para a Ciência e a Tecnologia, I.P., in the framework of the Project UIDB/00326/2025 and UIDP/00326/2025.